OpenStudy (anonymous):

Solve for x: –4(3x – 2) = 6x + 2

OpenStudy (anonymous):

so far i got -1

OpenStudy (anonymous):

-12x+8 = 6x+2 -12x-6x = 2-8 -18x = -6 18x = 6 x = 1/3

OpenStudy (anonymous):

k